Iron Man is transported by one of his foes to the planet Magma, where he is forced to participate in gladiatorial games with the native Magmans. While no one is severely injured during these battles, winners do get an extra piece of Magman blueberry pie for dessert.

The size of Iron Man's opponents is determined by alleles at five loci in the Magman genome: we shall label them A1, a1, A2, a2, ..., A5, a5. Each capital allele is an increaser allele, adding the same amount to the size and weight (musculature) of the Magman competitor, and each lower-case allele is a null allele, adding nothing. Iron Man knows that the smallest opponent he has faced, who is homozygous for all lower-case alleles, has a Magman Musculature Index (MMI) of 25.

Iron Man also knows that his 2nd opponent, with a multi-locus genotype of: A1a1 a2a2 a3a3 A4a4 a5a5, has a MMI of 85.

Given this information, what is the MMI of the strongest opponent Iron Man must face; that is, the opponent who is homozygous for all capital alleles at all five loci?

A. 400.

B. 330.

C. The answer cannot be determined.

D. 300.

E. 325.
Answer #1

we already know basal MMI when all alleles are recessive(a1a1a2a2a3a3a4a4a5a5) is 25

presence of two dominant alleles (A1a1a2a2a3a3A4a4a5a5) resulted in increase of MMI to 85

in otherwords addition of two dominant alleles added an MMI of 60 over the basal MMI of 85

i.e each allele added 60/2 = 30 MMI

there will ten dominant alleles when all the 5 genes are homozygous dominant (A1A1A2A2A3A3A4A4A5A5)

therefore and addition of 10 x 30 = 300 MMI will occure over the basal MMI

therefore strongest person Iron man will encounter have an MMI = Basal MMI + Added MMI from dominant alleles =25+300 = 325

Ans. E. 325